OpenStudy (anonymous):

Annie, Ben and Claire are training for MATHCOUNTS. On January 1st, they each started to solve a large packet of problems that their coach gave them. Each of them worked at a constant rate. Annie solved six problems each day and finished four days after Ben did. Claire solved three more problems each day than Ben did and finished two days before he did. How many problems were in the packet their coach gave them?

OpenStudy (anonymous):

A:6d = x d = x/6 B: x/6 - 4 = B's number of days B's rate is x/(x/6 - 4 ) C's rate is x/(x/6 - 4 ) + 3 C's number of days = x/6 - 4 + 2 = x/6 -2

OpenStudy (anonymous):

C's number of days = x/6 - 4 - 2 = x/6 - 6

OpenStudy (anonymous):

(x/6 - 6) (x/(x/6 - 4 ) + 3) = x

OpenStudy (anonymous):

x must be 72

OpenStudy (anonymous):

i guess common denominator first

OpenStudy (anonymous):

math counts?

OpenStudy (anonymous):

There are probably shortcuts for this problem